### v2

# Alteraxial Phonons in Collinear Magnets

Abstract
Axial phonons possessing nonzero angular momentum and resulting magnetic moment can couple to magnetic order. The rich magnetic structures enable phonon angular momentum (PAM) to acquire momentum-space textures analogous to electronic spin structures. However, a systematic framework for classifying these textures, especially their potential higher-order multipolar patterns, has remained elusive. Here, by employing magnetic point group analysis, we develop a complete classification of phonons in collinear magnets, spanning ferro-, antiferro-, and alter-magnets. Our theory distinguishes four fundamental types of magnetic phonons based on symmetry and PAM wave-pattern parity. Strikingly, we reveal a full sequence of axial phonons exhibiting higher-order-wave (from $p$- to $j$-wave) PAM patterns covering both odd and even parities, which we term alteraxial phonons. Our high-throughput calculations predict hundreds of magnetic candidates hosting such alteraxial phonons. Our work establishes a symmetry-guided design principle for axial phonons and related phenomena in magnetic materials.
Introduction. Circularly polarized phonons with finite angular momentum [1, 2, 3], recently termed axial phonons [4], have attracted growing research interest over the past decade [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 11, 10, 18, 19, 20, 21]. Intriguingly, nonzero phonon angular momentum (PAM) can yield an effective phonon magnetic moment[22, 13, 23, 24, 25, 26, 27], which enables the coupling of axial phonons to both external magnetic fields and internal magnetic moments [28, 29, 30, 31, 32, 33, 34, 8, 35, 36, 37]. Such couplings have been shown to give rise to various phono-magnetic effect, including the phonon Zeeman effect [38, 39], the ultrafast Einstein-de Haas effect [40, 33] and the phonon Barnett effect [22, 32]. Specifically, the phonon Zeeman effect can induce observable frequency splittings between left- and right-circularly polarized phonons in magnetic materials, as demonstrated in Co3Sn2S2 [41, 42, 43]. Despite recent progress, the interplay between axial phonons and magnetic order still remains largely unexplored.
Figure 1: Schematics of three distinct types of circularly polarized phonons in collinear magnets. Circular lattice vibrations in real space (top row), $\Gamma$-centered phonon spectra (middle row) and angular momentum patterns in momentum space (bottom row) for (a) ferro-axial, (b) antiferro-nonaxial and (c) alter-axial phonons in collinear magnets, respectively. The magnetization direction is set as the quantization axis for the phonon angular momentum. Blue and red colors represent the left- and right-handed circular polarization, respectively.
Table 1: Complete classification of phonons in collinear magnets. Phonons are categorized into four types by magnetic point group symmetry, with distinct phonon angular momentum (PAM) wave patterns: nonaxial, centro-even-wave ($s$-, $d$-, $g$-,$i$-wave), noncentro-odd-wave ($p$-, $f$-, $h$-,$j$-wave), and noncentro-even-wave ($s$-, $d$-, $g$-,$i$-wave). The $s$-wave pattern describes ferro-axial phonons, whereas higher-order patterns ($p$- to $j$-wave) correspond to alteraxial phonons. Magnetic point groups are given in the UNI notation [44]. Parentheses are used to specify the orientation of the magnetic axis.
Type Wave Magnetic Point Group
Non-axial N/A
$1^{\prime}, 6/mmm6^{\prime}/mmm^{\prime}$, $6/m^{\prime}m^{\prime}m^{\prime}$, $m^{\prime}m^{\prime}m^{\prime}$, $1^{\prime}, 6/m6^{\prime}/m$, $^{\prime}mm, 6/m6/m^{\prime}$, $4/m^{\prime}mm$, $4/m^{\prime}m^{\prime}m^{\prime}$,
$4^{\prime}/m^{\prime}m^{\prime}m$, $4/m^{\prime}$, $4/mmm.1^{\prime}$, $1^{\prime}, 4/m4^{\prime}/m^{\prime}$, $2^{\prime}/m$, $2/m^{\prime}$, $2/m.1^{\prime}$, $^{\prime}mm, mm$$mm.1^{\prime}$, $\overline{1}^{\prime}$, $\overline{1}.1^{\prime}$, $\overline{3}^{\prime}m$,
$\overline{3}^{\prime}m^{\prime}$, $\overline{3}m.1^{\prime}$, $\overline{3}.1^{\prime}$, $\overline{3}^{\prime}$
Centro-even $s$ $6/m.1$, $m^{\prime}m^{\prime}m(z)$, $2/m.1(y)$, $4/mm^{\prime}m^{\prime}$, $\overline{3}m^{\prime}$, $4/m.1$, $\overline{3}.1$, $\overline{1}.1$, $2^{\prime}/m^{\prime}(xz)$, $6/mm^{\prime}m^{\prime}$
$d$ $2^{\prime}/m^{\prime}(y)$, $^{\prime}m^{\prime}m(y), mm$m.1mmm.1, $2/m.1(xz)$, $mmm.1$, $^{\prime}m^{\prime}m(x), mm$m.1mmm.1, $4^{\prime}/m$, $4^{\prime}/mm^{\prime}m$
$g$ $4/mmm.1$, $\overline{3}m.1$, $6^{\prime}/m^{\prime}mm^{\prime}$, $6^{\prime}/m^{\prime}$
$i$ $6/mmm.1$
Noncentro-odd $p$
$6.1^{\prime}$, $m.1^{\prime}$, $2.1^{\prime}$, $4^{\prime}$, $222.1^{\prime}$, $6^{\prime}22^{\prime}$, $622.1^{\prime}$, $422.1^{\prime}$, $32.1$, $622.1$, $222.1$, $32.1^{\prime}$, $2.1(xz)$, $4^{\prime}22^{\prime}$, $6^{\prime}$, $3.1^{\prime}$,
[…]
$f$ $\overline{4}.1^{\prime}$, $mm2.1^{\prime}(z)$, $\overline{4}2m.1$, $\overline{4}2m.1^{\prime}$, $3m.1$, $\overline{6}m2.1^{\prime}$, $\overline{6}^{\prime}m2^{\prime}$, $3m.1^{\prime}$, $\overline{6}^{\prime}m^{\prime}2$, $\overline{6}m2.1$, $\overline{6}^{\prime}$, $\overline{6}.1^{\prime}$
$h$ $4mm.1^{\prime}$
$j$ $6mm.1^{\prime}$
Noncentro-even $s$
$m^{\prime}m2^{\prime}(y)$, $2^{\prime}2^{\prime}2(z)$, $62^{\prime}2^{\prime}$, $2^{\prime}(xz)$, $1.1$, $6.1$, $m^{\prime}m^{\prime}2(z)$, $2.1(y)$, $3m^{\prime}$, $3.1$, $32^{\prime}$, $4m^{\prime}m^{\prime}$, $m^{\prime}(xz)$, $m.1(y)$,
$\overline{6}m^{\prime}2^{\prime}$, $6m^{\prime}m^{\prime}$, $\overline{6}.1$, $42^{\prime}2^{\prime}$, $4.1$, $\overline{4}2^{\prime}m^{\prime}$, $\overline{4}.1$
$d$ $\overline{4}^{\prime}$, $mm2.1(z)$, $\overline{4}^{\prime}2m^{\prime}$, $m^{\prime}m2^{\prime}(z)$, $\overline{4}^{\prime}2^{\prime}m$, $4^{\prime}m^{\prime}m$
$g$ $4mm.1$, $6^{\prime}mm^{\prime}$
$i$ $6mm.1$
A key open question regarding axial phonons in magnetic materials is how crystal and magnetic structures shape momentum-space phonon Zeeman effects and resultant PAM patterns. Analogous to the classification of collinear magnets (ferro-, antiferro-, alter-magnets) by their magnetic structures [45, 46, 47], phonons in collinear magnets can, in principle, be similarly categorized into such distinct types, as schematized in Fig. 1. Ferro-axial phonons [Fig. 1(a)] are characterized by circular lattice vibrations of a single handedness in real space and circularly polarized phonon spectra in momentum space. Antiferro-nonaxial phonons [Fig. 1(b)] feature $\mathcal{PT}$-symmetric circular vibrations in real space and unpolarized phonon spectra with vanishing PAM in momentum space. Alteraxial phonons [Fig. 1(c)] exhibit real-space alternating circular vibrations that are not connected by $\mathcal{PT}$ symmetry and alternating circularly polarized phonon spectra with wave-modulated PAM patterns in momentum space. However, a comprehensive group-theoretical study of phonons in collinear magnets is still lacking, thereby significantly hindering the identification of magnetic materials hosting alteraxial phonons.
Figure 2: High-throughput phonon calculations in collinear magnets. The workflow (a) consists of three parts: The gray-shaded part filters out collinear magnets from material databases and provides their crystal and magnetic structures along with the corresponding magnetic point groups (MPG); The cyan-shaded part classifies the phonon type according to the MPG, where the MPG integer denotes the number of symmetry-enforced nodal surfaces of vanishing phonon angular momentum crossing the $\Gamma$ point; The pink-shaded part performs ab initio calculations of magnetic phonon spectra. (b, c) High-throughput screening for all 1029 collinear magnets from the MAGNDATA database, with (b) distribution of materials across the four distinct phonon types and (c) number of materials hosting different wave patterns within the three axial phonon families.
In this work, based on magnetic point group (MPG) analysis, we perform a complete group-theoretical classification of phonons in collinear magnets. Specifically, we divide them into four categories distinguished by symmetry and momentum-space PAM wave patterns: one nonaxial type and three axial types, comprising centro-even-, noncentro-odd-, and noncentro-even-wave phonons. Remarkably, beyond the ferro-axial ($s$-wave) phonon, we reveal higher-order-wave PAM patterns featuring one to seven nodal surfaces of vanishing PAM crossing the $\Gamma$ point. These correspond to the complete series from $p$- to $j$-wave alteraxial phonons, encompassing both odd- and even-parity wave patterns [see Supplemental Material (SM) [48]]. We have also performed high-throughput calculations of magnetic materials, and predicted hundreds of candidates hosting alteraxial phonons. We further present ab initio magnetic phonon spectra and PAM patterns for three representative alteraxial phonon materials, namely, CrSb, Cr2SbAs, and MnSe, spanning distinct categories. By establishing a systematic framework for axial phonons in collinear magnets, our work paves the way for phono-magnetic effects.
Symmetry-based classification scheme. The concept of alteraxial phonons in collinear magnets is based on the magnitude of the PAM defined with the magnetization direction (labeled as $\vec{e}_{\mathrm{m}}$ direction) as the quantization axis. For a single phonon mode $\epsilon_{\bf{q}}$ with $\bf{q}$ denoting its wavevector, this magnitude corresponds to the phonon circular polarization along $\vec{e}_{\mathrm{m}}$ direction, which is given by
$j_{\textrm{m}}(\bf{q})=\langle\epsilon_{\bf{q}}|\ \big(|\textrm{R}_{\textrm{m}}\rangle\langle\textrm{R}_{\textrm{m}}|-|\textrm{L}_{\textrm{m}}\rangle\langle\textrm{L}_{\textrm{m}}|\big)\ |\epsilon_{\bf{q}}\rangle\hbar.$ (1)
Here, $|\epsilon_{\bf{q}}\rangle$ denotes the phonon-mode eigenvector, $\hbar$ is the Planck’s constant, and $|\textrm{R}_{\textrm{m}}\rangle$ ($|\textrm{L}_{\textrm{m}}\rangle$) represents the right-handed (left-handed) circularly polarized eigenstate with angular momentum $\hbar$ (-$\hbar$) along $\vec{e}_{\mathrm{m}}$ direction.
Throughout this work, we consider $\Gamma$-centered phonon modes in collinear magnets, whose symmetry properties are characterized by a MPG denoted as $G$ [44]. Under a symmetry operation $\mathcal{G}\in G$, the projected PAM $j_{\mathrm{m}}(\bf{q})$ satisfies
$j_{\textrm{m}}(\mathcal{G}\textbf{q})=\eta(\mathcal{T})\chi(\mathcal{R})j_{\textrm{m}}(\textbf{q}).$ (2)
Here, $\mathcal{T}$ is the time-reversal operation, and $\eta(\mathcal{T})=-1$ (+1) when $\mathcal{T}$ operation is (not) contained in $\mathcal{G}$. $\mathcal{R}$ denotes the point-group part operation in $\mathcal{G}$, and $\chi(\mathcal{R})=+1$ (1)(-1) if $\mathcal{R}$ preserves (switches) the handedness of the circularly polarized states $|\textrm{R}_{\textrm{m}}\rangle$ and $|\textrm{L}_{\textrm{m}}\rangle$. Specifically, time-reversal symmetry $\mathcal{T}$ yields $j_{\textrm{m}}(-\textbf{q})=-j_{\textrm{m}}(\textbf{q})$, while inversion symmetry $\mathcal{P}$ gives $j_{\textrm{m}}(-\textbf{q})=j_{\textrm{m}}(\textbf{q})$. Consequently, in the presence of the combined $\mathcal{PT}$ symmetry, $j_{\textrm{m}}(\textbf{q})=0$ is always satisfied, and the corresponding phonons are nonaxial.
Breaking the combined $\mathcal{PT}$ symmetry in collinear magnets enables a nonzero $j_{\textrm{m}}(\textbf{q})$, thereby leading to the emergence of axial phonon modes. Interestingly, if at least one symmetry operation yields $\eta(\mathcal{T})\chi(\mathcal{R})=-1$, then $j_{\textrm{m}}$ vanishes at the $\Gamma$ point, while alternating in sign across the surrounding Brillouin zone (BZ). This results in nodal surfaces of vanishing $j_{\textrm{m}}(\textbf{q})$ which pass through the $\Gamma$ point in the three-dimensional BZ. The number of such nodal surfaces, determined from MPG constraints in Eq. (2), ranges from 1 to 7, corresponding to $p$-, $d$- , $f$- , $g$- , $h$- , $i$- , $j$-wave alteraxial phonons, respectively (see the SM for detailed illustrations). In contrast, if no symmetry operations gives $\eta(\mathcal{T})\chi(\mathcal{R})=-1$, then no such symmetry-enforced nodal surfaces exist, and $j_{\textrm{m}}(\textbf{q})$ generically exhibits $s$-wave feature in the BZ, thereby termed ferro-axial phonons.
Based on the above symmetry analysis, we categorize phonons in collinear magnets into four distinct types according to the underlying symmetry and the wave parity of the PAM pattern: (i) $\mathcal{PT}$-symmetric nonaxial phonons; (ii) Centro-even-wave ($s$-, $d$-, $g$-, $i$-wave) axial phonons; (iii) Noncentro-odd-wave ($p$-, $f$-, $h$-, $j$-wave) axial phonons; (iv) Noncentro-even-wave ($s$-, $d$-, $g$-, $i$-wave) axial phonons. The complete classification for all collinear MPGs is presented in Table I. Notably, since $j_{\textrm{m}}(\textbf{q})$ is odd under time-reversal symmetry $\mathcal{T}$, the breaking of $\mathcal{T}$ is a necessary condition for the emergence of both type (ii) and type (iv) phonons which feature even-wave PAM and are inherent to magnetic systems. In magnetic materials, such $\mathcal{T}$-breaking terms typically arise from spin-phonon [51, 31, 52, 35, 39] or magnon-phonon couplings [53, 54, 55, 56, 57, 58, 59, 60, 61], which usually constitute perturbative contributions.
Figure 3: Phonon spectra and phonon angular momentum (PAM) patterns from ab initio calcualtions on representative materials hosting three different types of alteraxial phonons. CrSb (centro-even-wave), Cr2SbAs (noncentro-odd-wave), and MnSe (noncentro-even-wave) are selected as illustrative cases. For each material, the top row shows the crystal and magnetic structures, the middle (bottom) row presents the $\mathcal{T}$-symmetric ($\mathcal{T}$-broken) phonon spectra and $\Gamma$-centered PAM patterns in the absence (presence) of $\mathcal{T}$-breaking perturbations.
High-throughput calculations. To search for realistic collinear magnets hosting alteraxial phonons, we performed high-throughput calculations on magnetic materials from databases such as MAGNDATA [62, 63] and Materials Project [64]. The calculation workflow, as depicted in Fig. 2(a), comprises two primary tasks after obtaining the crystal and magnetic structures, namely, determining the MPG-based phonon type and conducting ab initio calcualtions of magnetic phonon spectra. It is worth mentioning that our calcualtions incorporate the Raman-type spin-phonon coupling induced from electron-phonon interaction as the $\mathcal{T}$-breaking perturbation [65], which is implemented via a recently developed method by some of the authors [39] (see the SM for details).
Figures 2(b) and 2(c) summarized our results for all 1029 collinear magnets from the MAGNDATA database [see SM [48] for the results of Material Project database]. Of these 1,029 materials, 619 (60.2%) host nonaxial phonons. The remaining 410 materials host axial phonons, distributed as follows: 271 (26.3%) are centro-even-wave, 93 (9.0%) are noncentro-odd-wave, and 46 (4.5%) are noncentro-even-wave. Notably, we uncover 237 magnets that host alteraxial phonons containing both odd- and even-parity wave patterns, in addition to 173 materials exhibiting nodeless $s$-wave axial phonons.
Representative materials. Our calculations reveal that alteraxial phonons featuring higher-order PAM wave patterns exist in collinear magnets across all magnetic types, including ferro-, antiferro-, and alter-magnets. Inspired by recent progresses in altermagnetism, we select three altermagnets as representative materials (see the SM for other magnets), each hosting a distinct type of alteraxial phonon: CrSb [46] (centro-even-wave); Cr2SbAs (noncentro-odd-wave); and MnSe (noncentro-even-wave), as detailed below.
CrSb exhibits an out-of-plane collinear magnetic order with antiparallel-aligned moments on Cr atoms [see Fig.3(a)], which belongs to the MPG $6^{\prime}/m^{\prime}mm^{\prime}$. Within the spin-group description, it is classified as a $g$-wave altermagnet [46]. It possesses inversion symmetry with Cr atoms acting as the inversion center. Consequently, in the absence of $\mathcal{T}$-breaking perturbation, the phonon spetrum of CrSb preserves the combined $\mathcal{PT}$ symmetry, yielding nonaxial phonons with vanishing PAM [Fig. 3(d)]. Notably, when $T$-breaking perturbation is incorparated, alteraxial phonons with even parity $g$-wave PAM pattern emerge. This is evidenced by the phonon spectrum and PAM distribution [Fig. 3(g)] characterized by four nodal-surfaces with vanishing PAM crossing the $\Gamma$ point (three vertical surfaces connected by $C_{3z}$ and one horizontal $q_{z}=0$ surface enforced by $C_{2z}\mathcal{T}$ symmetry).
Cr2SbAs is obtained by substituting As for one Sb site from CrSb [see Fig. 3(b)], thereby breaking the inversion symmetry and lowering the MPG to $\overline{6}^{\prime}m^{\prime}2$. Without the $\mathcal{T}$-breaking perturbation, the phonon spectrum exhibits odd-parity $f$-wave PAM pattern, with three vertical zero-PAM nodal surfaces passing through the $\Gamma$ point [Fig. 3(e)]. This $f$-wave pattern persists even when the $\mathcal{T}$-breaking perturbation is present, as shown in Fig. 3(h). In fact, the $f$-wave pattern results from the removal of the $q_{z}=0$ nodal surface in CrSb due to the broken $C_{2z}\mathcal{T}$ symmetry. Interestingly, Cr2SbAs remains a $g$-wave altermagnet, since its electronic structure is described by spin groups [66, 67] distinct from the MPG description of its phonon spectrum.
MnSe is a noncentrosymmetric magnet hosting out-of-plane antiparallel alignment of Mn moments [see Fig. 3(c)], described by the MPG $6^{\prime}mm^{\prime}$. As shown in Fig. 3(f), the unperturbed phonon spectrum of MnSe with preserved $\mathcal{T}$ symmetry exhibits an odd-parity $j$-wave PAM pattern, featuring seven zero-PAM nodal surfaces crossing the $\Gamma$ point: six vertical surfaces related by $C_{6z}$ symmetry, and one horizontal surface ($q_{z}=0$) enforced by $C_{2z}\mathcal{T}$ symmetry. Under a $\mathcal{T}$-breaking perturbation, the $C_{6z}$ symmetry is reduced to $C_{6z}\mathcal{T}$, whereas $C_{2z}\mathcal{T}$ remains. This shifts three vertical nodal surfaces from the $\Gamma$ point, which in turn changes the $\Gamma$-centered PAM pattern from odd-parity $j$-wave to even parity $g$-wave. It is worth noting that all noncentro-even-wave PAM patterns can be derived from an unperturbed higher-order-wave odd-parity pattern that preserves $\mathcal{T}$ symmetry, and the full derivation based on compatibility relations for each case is detailed in the SM.
Conclusion. We have established a comprehensive group-theoretical framework to systematically classify axial phonons in collinear magnets and uncovered a complete hierarchy of alteraxial phonons characterized by higher-order-wave PAM patterns. Our high-throughput screening identifies numerous candidate materials and ab initio calculations are further performed on representative materials to confirm alteraxial phonons as well as their PAM patterns. Our work not only significantly enriches the family of axial phonons, but also opens a new avenue in phonon engineering via magneto-phononic effects.
Acknowledgments. This work is partly supported by National Key Projects for Research and Development of China (Grants No. 2024YFA1409100, No. 2021YFA1400400, No. 2022YFA1403602, and No. 2023YFA1407001), the Natural Science Foundation of Jiangsu Province (Grants No. BK20253012, No. BK20252117, No. BK20233001, No. BK20243011, and No. BK20220032), the Natural Science Foundation of China (Grants No. 12534007, No. 92365203 and No.12174182) and the e-Science Center of Collaborative Innovation Center of Advanced Microstructures.